# Study of charged Lepton Flavor Violation in electron muon interactions

### v1

@@ Abstract
With the improvement of muon acceleration technology, it has received great interest to exploit high-energy muon beams for collision or target experiments. We investigate possible charged Lepton Flavor Violation (cLFV) processes mediated by an extra massive neutral gauge boson $Z^{\prime}$ in electron muon interactions, either at a proposed electron muon collider or in a fixed target experiment with high-energy muon beam hitting electrons in the target. Based on Monte Carlo calculations and fast detector simulations, we study in detail our signal and possible backgrounds, giving the sensitivity results of cLFV signals at the 90% confidence level. Compared with current and prospective limits set by other experiments, electron muon interactions demonstrate significant advantages in the cLFV coupling strength sensitivity with $\tau$ in the final states. In addition, a special cLFV coupling combination, $\lambda_{e\mu}\times\lambda_{\mu\mu}$, can also be probed in our proposal.
@@ 1 Introduction
In this study we also investigate a fixed electron-target experiment with a muon beam in addition to the $e\mu$ collider. Since lepton flavor in the initial state is non-zero, electron muon interaction can strongly avoid many potential background processes which would occur at different-sign muon colliders or electron-positron colliders, thus possessing higher sensitivity to new physics signals, typically the charged Lepton Flavor Violation (cLFV) processes.
In the SM framework, the cLFV processes are strongly suppressed due to the tiny mass of neutrinos, hence unobservable in the current experiments yet. However, it may be much enhanced in various BSM models, such as super-symmetry (SUSY) Barbier:2004ez , leptoquark Dorsner:2016wpm , two-Higgs-doublet Branco:2011iw , and the heavy neutral gauge bosons $Z^{\prime}$ Langacker:2008yv studied in this paper. In the past decades, searches for the cLFV process were performed in different channels with several approaches, typically the high intensity muon-based experiments including $\mu^{+}\to e^{+}\gamma$ (MEG) MEG:2016leq , $\mu^{+}\to e^{+}e^{+}e^{-}$ (SINDRUM) SINDRUM:1987nra and $\mu^{-}N\to e^{-}N$ (SINDRUM \@slowromancapii@) SINDRUMII:2006dvw ; SINDRUMII:1993gxf ; SINDRUMII:1996fti ; SINDRUMII:1998mwd , as well as the collider-based searches for cLFV decays of $Z$ ATLAS:2014vur ; OPAL:1995grn ; DELPHI:1996iox , Higgs ATLAS:2019pmk ; CMS:2017con and several hadron resonances KTeV:2007cvy ; CLEO:2008lxu ; BaBar:2021loj ; BESIII:2022exh . Meanwhile, there will be continuous new experiments conducted in the near future to constantly improve the existing limits, such as MEG\@slowromancapii@ MEGII:2018kmf , Mu3e Mu3e:2020gyw , COMET COMET:2018auw and Mu2e Mu2e:2014fns .
@@ 2 Physics processes and Monte Carlo simulation
In this study, $Z^{\prime}$ is considered to have the same coupling and chiral structure as the standard model $Z^{0}$, but allows for the lepton flavor violation, similarly as Ref. ATLAS:2018mrn ; Li:2023lin . The coupling strength of the $Z^{\prime}$ and leptons can be described by a matrix $\lambda$ as Eq. 1
@@
Generally, it represents the strength of the cLFV couplings relative to the SM couplings, assuming that the diagonal elements are 1, while the off-diagonal elements are usually a higher order of magnitude. Therefore, the cases that break lepton flavor conservation twice would not be considered in the further study. After introducing the $Z^{\prime}$ boson, the cLFV processes mentioned in Sec.1 can be enhanced by the diagrams shown in Fig.1. The branching ratio limits would be transformed to the coupling $\lambda_{ij}$ Bernabeu:1993ta ; Langacker:2000ju and compared with our results based on $e\mu$ interaction.
@@ 2.2 Physics processes
The signal cLFV processes studied in this paper are listed in Tab.2. There are two diagrams for the $Z^{\prime}$ mediated cLFV process $\mu^{+}e^{-}\to l^{+}l^{-}$, as shown in Fig.2 (taking $\mu^{+}e^{-}\to e^{+}e^{-}$ as an example). In particular, the s-channel is not included in the processes $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$. In the simulation, all coupling strengths are considered as 1. And the mass of $Z^{\prime}$ floats from 0.2 GeV to 5 TeV in the electron muon collision experiment, and within 0.50 GeV in the electron-target experiment with a muon beam.
@@
While several background processes may occur on the collider and affect the signal that we are interested in, these processes include the standard model backgrounds divided by the number of final state particles and accidental background caused by particle mis-identification. In conclusion, the specific signals and their background processes are shown in Tab.2.
@@ 2.3 Event generation and simulation
Both signal and background events are simulated by MadGraph5_aMC@NLO(MG) Madgraph version 3.1.1, which is one of the key tools for Monte Carlo event generation in high energy physics, then showered and hadronized by Pythia8 Pythia . Next, Delphes Delphes version 3.5.1 is utilized to simulate detector effects with the default configuration card for the detector at the muon collider.
In the Monte Carlo generation, some preliminary requirements are applied to remove the physically unreasonable events. In $e\mu$ collisions, the transverse momentum of charged leptons is required to satisfy $p_{T}>$ 10 GeV and the absolute pseudo-rapidity of charged leptons $\lvert\eta\rvert>$ 2.5. While in the muon-beam electron-target simulation, the filtering criteria of $p_{T}$ and $\lvert\eta\rvert$ would be relaxed. Then in detector simulation, parameters such as efficiency of particle detection are set according to the Delphes cards.
There is some difference between the $\tau$ simulation and $e/\mu$. For those final states with $\tau$ in Tab.2, although they can go through any decay chains in reality, in this study we only consider the hadronic decay channels (about 60% of the total decay) and reconstruct it by the Jet information.
@@ 3.1.1 Background study
After setting the preliminary requirements as mentioned in Sec.2.3, signal candidate events should be with the same charged leptons corresponding to the signal in the final state. Since the initial flavor in $e\mu$ collision is non-zero, vast majority of the SM background are forbidden, while the remaining portion also exhibits significant kinematic differences from the signal processes, especially in $ee$ or $\mu\mu$ channel. The invariant mass distributions of the final state dileptons of these two channels are shown in Fig.LABEL:fig:mlldistribution1.
Only considering the interval near the center-of-mass energy, the SM background values are extremely low. And compared with a similar study on the different-signs election or muon collider Li:2023lin , $e\mu$ collision has a cleaner signal window. Due to the low level of the SM background, we also investigated the accidental background caused by $e\mu$ mis-identification, where a final state muon is assigned the mass of the electron, or vice versa. Typically it would let the $e\mu$ scattering process coming into the background of $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$. The probability of $e\mu$ mis-identification is set as 10-6. The invariant mass distribution of this process is extremely close to the signal.
while for $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$ since the reconstruction of $\tau$ would inevitably result in a certain loss of energy, there is a considerable overlap in the signal and background distributions. It will be optimized in the next section.
@@ 3.1.2 Sensitivity result
Based on the distributions of the signal and background, we truncate the invariant mass to remove the events with significantly deviating from center-of-mass energy. The specific truncation point in $\tau$ channels will be determined by scanning and selecting the maximum value of $S/(a/2+\sqrt{B})$, where $S$ is the number of signal processes, $B$ is the total number of the weighted background and $a$ is the significant value which is considered as 3 Punzi:2003bu . The weight is defined by $n_{x}=\sigma_{x}L/N$, where $\sigma_{x}$ is the cross section of each process, $L$ is the luminosity and $N$ is the generated number.
Then the binned histograms of leptons $p_{T}$ distributions are utilized for the statistic analysis. The test statistics $Z_{i}$ is calculated by $Z_{i}:=2[n_{i}-b_{i}+b_{i}\ln{(n_{i}/b_{i})}]$ for 90% confidence level (C.L.) exclusion, where $i$ is the index of each bin, $n$ is the weighted number of observed events including signal and background, and $b$ is the weighted number of background Cowan:2010js . Then the total $Z=\Sigma_{i}Z_{i}$ would be subject to a $\chi^{2}$ distribution. The degree of freedom is defined by the number of bins. By iteration, we can obtain the signal cross section of 90% C.L. exclusion in Fig.4.
Then based on MG calculation we can get the corresponding value of $\lambda_{ll^{\prime}}\times\lambda_{\rm SM}$, as shown in Fig.5. The $\lambda_{e\mu}$ and $\lambda_{e\tau}$ results are calculated at two energy points, and the luminosity is considered as 0.3 ab-1 or 5 ab-1. Several current limits and prospective limits are also included to compare with our results. The constraint of branching ratio from other experiments are concluded in Tab.3. It is sensible that the limits given by other experiments in $\tau$ channels are much more conservative than in $e\mu$ channels. But in our collider study, they may be similar since the main influencing factor here is the signal cross section while the cross sections of different signals are quite similar.
@@
Table 3: Summary of current and prospective limits from other experiments at 90% C.L.
@@
Figure 5: The 90% C.L. upper limit of $\lambda_{e\mu}\times\lambda_{\rm SM}$ (a) and $\lambda_{e\tau}\times\lambda_{\rm SM}$ (b). The curves are graphed with respect to $M_{Z^{{}^{\prime}}}$, representing the limits of the cross section times branching ratio. Additionally, exclusion lines from both present low-energy experiments (shown as solid lines) and future experiments (shown as dashed lines) are included in the plot.
@@
In the $\lambda_{e\mu}$ channel, the strictest constraint comes from the results of $\mu-e$ conversion, and our results have weaker competitiveness among those high intensity muon-based experiments. While in the $\lambda_{e\tau}$ channel, current limits of $\tau\to eee$ and $\tau\to e\mu\mu$ perform the best across the entire interval, but the orders of magnitude are much lower than the results of $\lambda_{e\mu}$. Compared with our results, the constraints of 1.55 TeV $e\mu$ collider are more stringent than other existing results when $M_{Z^{\prime}}>0.5$ TeV, and of 346 GeV collider it is about the entire $M_{Z^{\prime}}$ interval. Even compared with the prospective limits of those processes on Belle II, our results still have certain advantages around the resonance region.
@@
On the other side, comparing the results of $e\mu$ collision and $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ collisions Li:2023lin , based on the same luminosity, the results of $e\mu$ are better than $\mu^{+}\mu^{-}$ by an order of magnitude at their respective resonance points. But it cannot be ignored that the center-of-mass energy of $e\mu$ is relatively low, and the results of other experiments are more advantageous in the low-energy region. Not only that, for $e\mu$ collider it is difficult to let the luminosity reach the same level as $\mu^{+}\mu^{-}$ collider in reality. Therefore, the advantages of these two collision scenarios still need to be considered based on more practical factors.
@@
To be precise that what we are comparing in Fig.5(a) and Fig.5(b) is the coupling $\lambda_{e\mu}(\lambda_{e\tau})\times\lambda_{\rm SM}$. Then based on the assumption that $Z^{\prime}$ has the same coupling structures and strengths as the standard model $Z^{0}$ as mentioned in Sec.2.1, that is, $\lambda_{\rm SM}=1$, we can naturally obtain the estimation of $\lambda_{e\mu}$ or $\lambda_{e\tau}$. But more strictly speaking, this requirement may not exist in more universal models, and we need to consider the biases of standard model couplings. In this way, each process may give a product of different coupling strength, for example $\lambda_{e\mu}\times\lambda_{ee}$ ($\mu\to eee$), and $\lambda_{e\mu}\times\lambda_{ll}$ ($\mu\to e\gamma$) where $l$ represent $e$ or $\mu$. While for the high intensity muon-based experiments, they were unable to measure the coupling $\lambda_{e\mu}\times\lambda_{\mu\mu}$. This coupling requires the existence of a $Z^{\prime}-e-\mu$ vertex and a $Z^{\prime}-\mu-\mu$ vertex, which is almost impossible to achieve except for the direct interactions of leptons. On the $e\mu$ collider, we can obtain the measurement of this coupling through $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$, as shown in Fig.6, indicating that it has similar sensitivity as $\mu^{+}e^{-}\to e^{+}e^{-}$.
@@
Figure 6: Sensitivity of the coupling strength $\lambda_{\mu\mu}\times\lambda_{e\mu}$ from $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ channel.

### v3

@@ Abstract
With the improvement of muon acceleration technology, it has received great interests to exploit high-energy muon beams for collision or target experiments. We investigate the possible charged Lepton Flavor Violation (cLFV) processes mediated by an extra massive neutral gauge boson $Z^{\prime}$ in electron muon interactions, either at a proposed electron muon collider or in a fixed target experiment with high-energy muon beam hitting electrons in the target. Based on Monte Carlo calculations and fast detector simulations, we study in detail the signal and possible backgrounds, giving the sensitivity results of cLFV signals at the 90% confidence level. Compared with the current and prospective limits set by other experiments, electron muon interactions demonstrate significant advantages in the cLFV coupling strength sensitivity with $\tau$ in the final states. In addition, a special cLFV coupling combination, $\lambda_{e\mu}\times\lambda_{\mu\mu}$, can also be probed in our proposal.
@@ 1 Introduction
In this study we also investigate a fixed electron-target experiment with a muon beam in addition to the $e\mu$ collider. As lepton flavor in the initial state is non-zero, electron muon interaction can strongly avoid many potential background processes which would occur at different-sign muon colliders or electron-positron colliders, thus possessing higher sensitivity to new physics signals, typically the charged Lepton Flavor Violation (cLFV) processes.
In the SM framework, the cLFV processes are strongly suppressed due to the tiny mass of neutrinos, hence unobservable in the current experiments yet. However, it may be much enhanced in various BSM models, such as super-symmetry (SUSY) Barbier:2004ez , leptoquark Dorsner:2016wpm , two-Higgs-doublet Branco:2011iw , R-parity-violating (RPV) Minimal Super-symmetric Standard Model (MSSM) Choudhury:1996ia ; Chemtob:2004xr ; Cai:2024fkl , and the heavy neutral gauge boson $Z^{\prime}$ Langacker:2008yv studied in this paper. In the past decades, searches for the cLFV process were performed in different channels with several approaches, typically the high intensity muon-based experiments including $\mu^{+}\to e^{+}\gamma$ (MEG) MEG:2016leq , $\mu^{+}\to e^{+}e^{+}e^{-}$ (SINDRUM) SINDRUM:1987nra and $\mu^{-}N\to e^{-}N$ (SINDRUM \@slowromancapii@) SINDRUMII:2006dvw ; SINDRUMII:1993gxf ; SINDRUMII:1996fti ; SINDRUMII:1998mwd , as well as the collider-based searches for cLFV decays of $Z$ ATLAS:2014vur ; OPAL:1995grn ; DELPHI:1996iox , Higgs ATLAS:2019pmk ; CMS:2017con and several hadron resonances KTeV:2007cvy ; CLEO:2008lxu ; BaBar:2021loj ; BESIII:2022exh . Meanwhile, there will be continuous new experiments conducted in the near future to constantly improve the existing limits, such as MEG\@slowromancapii@ MEGII:2018kmf , Mu3e Mu3e:2020gyw , COMET COMET:2018auw and Mu2e Mu2e:2014fns .
@@ 2 Physics processes and Monte Carlo simulation
In this study, we consider that there are not $Z-Z^{\prime}$ mixing, and the extra $Z^{\prime}$ current has the same gauge coupling and chiral strength as the standard model $Z^{0}$, but allows for the lepton flavor violation, similarly as Ref. ATLAS:2018mrn ; Li:2023lin . The coupling strength of the $Z^{\prime}$ and leptons can be described by a matrix $\lambda$ as Eq. 1
@@
Generally, it represents the strength of the cLFV couplings relative to the SM couplings, assuming that the diagonal elements are 1, while the off-diagonal elements are usually a higher order of magnitude. Therefore, the cases that break lepton flavor conservation twice would not be considered in the further study. After introducing the $Z^{\prime}$ boson, the cLFV processes mentioned in Sec. 1 can be enhanced by the diagrams shown in Fig. 1. The branching ratio limits would be transformed to the coupling $\lambda_{ij}$ Langacker:2000ju and compared with our results based on $e\mu$ interaction.
@@ 2.2 Physics processes
The signal cLFV processes studied in this paper are listed in Tab. 2. There are two diagrams for the $Z^{\prime}$ mediated cLFV process $\mu^{+}e^{-}\to l^{+}l^{-}$, as shown in Fig. 2 (taking $\mu^{+}e^{-}\to e^{+}e^{-}$ as an example). In particular, the s-channel is not included in the processes $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$. In the simulation, all coupling strengths are considered as 1. And the mass of $Z^{\prime}$ floats from 0.2 TeV to 5 TeV in the electron muon collision experiment, and within 0.50 GeV in the electron-target experiment with a muon beam.
@@
While several background processes may occur on the collider and affect the signal that we are interested in, these processes include the standard model backgrounds divided by the number of final state particles and accidental background caused by particle mis-identification. Due to the suppression of phase space, we only consider the SM background with no more than 6 final state particles. In conclusion, the specific signals and their background processes are shown in Tab.2.
@@ 2.3 Event generation and simulation
Cross section and Monte Carlo events for each signal and background are calculated and simulated by MadGraph5_aMC@NLO(MG) Madgraph version 3.1.1, then showered and hadronized by Pythia8 Pythia . Next, Delphes Delphes version 3.5.1 is utilized to simulate detector effects with the default configuration card for the detector at the muon collider.
In the Monte Carlo generation, some preliminary requirements are applied to remove the physically unreasonable events. In $e\mu$ collisions, the transverse momentum of charged leptons is required to satisfy $p_{T}>$ 10 GeV and the absolute pseudo-rapidity of charged leptons $\lvert\eta\rvert>$ 2.5. While in the muon-beam electron-target simulation, the filtering criteria of $p_{T}$ and $\lvert\eta\rvert$ would be relaxed. Then in the detector simulation, the parameters such as particle detection efficiency are set according to the default detector configuration in the Delphes cards.
There are some differences between the $\tau$ simulation and $e/\mu$. For those final states with $\tau$ in Tab.2, although they can go through any decay chains in reality, in this study we only consider the hadronic decay channels (about 60% of the total decay) and reconstruct it by the JET information.
@@ 3.1.1 Background study
Unlike $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ collisions, since the initial lepton flavor in $e\mu$ collision is non-zero, vast majority of the SM backgrounds are forbidden, for example, $l^{+}l^{-}\to W^{+}W^{-}$ and $l^{+}l^{-}\to\tau^{+}\tau^{-}$. The typical diagrams for the SM background $\mu^{+}e^{-}\to l^{+}l^{-}\nu\bar{\nu}$ are shown in Fig.3.
Figure 3: Typical diagrams for the SM background $\mu^{+}e^{-}\to l^{+}l^{-}\nu\bar{\nu}$, taking $\mu^{+}e^{-}\to e^{+}e^{-}\nu\bar{\nu}$ as an example. Especially, the diagram (b) does not appear in $\mu^{+}e^{-}\to\mu^{+}\tau^{-}\nu\bar{\nu}$.
After setting the preliminary requirements as mentioned in Sec.2.3, the signal candidate events should have the same charged leptons corresponding to the signal in the final state. According to the invariant mass distributions of the final state di-leptons as shown in Fig. LABEL:fig:mlldistribution1, the remaining portion also exhibits significant kinematic differences from the signal processes, especially in $\mu^{+}e^{-}\to e^{+}e^{-}$ or $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ channel.
Only considering the interval near the center-of-mass energy, the SM background values are extremely low. And compared with a similar study on the different-signs electron or muon collider Li:2023lin , $e\mu$ collision has a cleaner signal window. Due to the low level of the SM background, we also investigated the accidental background caused by $e\mu$ mis-identification, where a final state muon is mis-identified as an electron, or vice versa. Typically it would let the $e\mu$ scattering process coming into the background of $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$. The probability of $e\mu$ mis-identification is set as 10-6. The invariant mass distribution of this process is extremely close to the signal.
While for $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$, since the reconstruction of $\tau$ would inevitably result in a certain loss of energy, there is a considerable overlap in the signal and background distributions. The analysis will be optimized in the next section.
@@ 3.1.2 Sensitivity result
Based on the distributions of the signal and background as shown in Fig. LABEL:fig:mlldistribution1, we truncate the invariant mass to remove the events with the reconstructed invariant mass significantly deviating from center-of-mass energy. For the cLFV channels without $\tau$, the truncation points are set as 1.4 TeV for the 1.55 TeV collider, and 300 GeV for the $\mu$TRISTAN. While in $\tau$ channels, it will be determined by scanning and selecting the maximum figure of merit (FOM) value of $S/(a/2+\sqrt{B})$, where $S$ is the number of signal events, $B$ is the weighted number of the background $\mu^{+}e^{-}\to\mu^{+}\tau^{-}\nu_{e}\bar{\nu}_{\tau}$ and $a$ is the significance value, which is usually set as 3 for new physics discoveries Punzi:2003bu . The weight for each process is defined by $n_{x}=\sigma_{x}L/N$, where $\sigma_{x}$ is the cross section, $L$ is the luminosity and $N$ is the generated event number, which is 200,000 for each process. When $L=0.3$ ab-1, the truncation point results are 1.27 TeV and 280 GeV for the two collider.
Then the binned histograms of leptons $p_{T}$ distributions are utilized for the statistical analysis. The test statistics $Z_{i}$ is calculated by $Z_{i}:=2[n_{i}-b_{i}+b_{i}\ln{(n_{i}/b_{i})}]$ for 90% confidence level (C.L.) exclusion, where $i$ is the index of each bin, $n$ is the weighted number of observed events including signal and background, and $b$ is the weighted number of background Cowan:2010js . Then the total $Z=\Sigma_{i}Z_{i}$ would be subject to a $\chi^{2}$ distribution. The degree of freedom is defined by the number of bins. By iteration, we can obtain the signal cross section of 90% C.L. exclusion, as shown in Fig. 5.
@@
Then based on MadGraph calculation we can get the corresponding value of $\lambda_{ll^{\prime}}\times\lambda_{\rm SM}$, as shown in Fig. 6. Currently the coupling $\lambda_{ij}$ are considered as positive real. Then the $\lambda_{e\mu}$ and $\lambda_{e\tau}$ results are calculated at two energy points, and the luminosity is 0.3 ab-1 or 5 ab-1. Several current limits and prospective limits are also included to compare with our results. The constraints of branching ratio from other experiments are concluded in Tab. 3. It is sensible that the limits given by other experiments in $\tau$ channels are much more conservative than in $e\mu$ channels. But in our collider study, they may be similar since the main influencing factor here is the signal cross section while the cross sections of different signals are quite similar.
Table 3: Summary of the current and prospective limits from other experiments at 90% C.L.
@@
Figure 6: The 90% C.L. upper limit of $\lambda_{e\mu}\times\lambda_{\rm SM}$ (a) and $\lambda_{e\tau}\times\lambda_{\rm SM}$ (b). The curves are graphed with respect to $M_{Z^{{}^{\prime}}}$, representing the limits of the cross section times branching ratio. Additionally, exclusion lines from both present low-energy experiments (shown as solid lines in purple, green and brown) and future experiments (shown as dashed lines in purple, green and brown) are included in the plot.
@@
In the $\lambda_{e\mu}$ channel, the strictest constraint comes from the results of $\mu-e$ conversion, and our results have weaker competitiveness among those high intensity muon-based experiments. While in the $\lambda_{e\tau}$ channel, current limits of $\tau\to eee$ and $\tau\to e\mu\mu$ perform the best across the entire interval, but the orders of magnitude are much lower than the results of $\lambda_{e\mu}$. Compared with our results, the constraints of 1.55 TeV $e\mu$ collider are more stringent than other existing results when $M_{Z^{\prime}}>0.5$ TeV, and for the 346 GeV collider it is about the entire $M_{Z^{\prime}}$ interval. Even compared with the prospective limits of those processes on Belle II, our results still have certain advantages around the resonance region.
@@
On the other side, comparing the results of $e\mu$ collision and $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ collisions Li:2023lin based on the same luminosity, the results of $e\mu$ are better than $\mu^{+}\mu^{-}$ by an order of magnitude at their respective resonance points. But it cannot be ignored that the center-of-mass energy of $e\mu$ is relatively low, and the results of other experiments are more advantageous in the low-energy region. Furthermore, it is difficult for $e\mu$ collider to reach the same luminosity as that of $\mu^{+}\mu^{-}$ collider in reality. Therefore, the advantages of these two collision scenarios still need to be considered based on more practical factors.
@@
To be precise, what we are comparing in Fig. 6(a) and Fig. 6(b) is the coupling $\lambda_{e\mu}(\lambda_{e\tau})\times\lambda_{\rm SM}$. Then based on the assumption that $Z^{\prime}$ has the same coupling structures and strengths as those of the standard model $Z^{0}$ mentioned in Sec. 2.1, which is $\lambda_{\rm SM}=1$, we can naturally obtain the estimation of $\lambda_{e\mu}$ or $\lambda_{e\tau}$. However, strictly speaking, this requirement may not exist in more universal models, and we need to consider the biases of standard model couplings. In this way, each process may give a product of different coupling strength, for example $\lambda_{e\mu}\times\lambda_{ee}$ ($\mu\to eee$) and $\lambda_{e\mu}\times\lambda_{ll}$ ($\mu\to e\gamma$), where $l$ represents $e$ or $\mu$. Specially, we can obtain the measurement of the coupling $\lambda_{e\mu}\times\lambda_{\mu\mu}$ through $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ based on the $e\mu$ collider, as shown in Fig. 7, indicating that it has similar sensitivity as $\mu^{+}e^{-}\to e^{+}e^{-}$. While for the current high intensity experiments, this coupling only appears in $\mu\to e\gamma$ together with $\lambda_{e\mu}\times\lambda_{ee}$ according to Ref. Langacker:2000ju , as
$\Gamma(\mu\to e\gamma)=\frac{\alpha G_{F}^{2}m^{3}_{\mu}M^{4}_{Z}}{4\pi^{4}M^{4}_{Z^{\prime}}}\left[\sin^{2}\theta_{W}\left(\sin^{2}\theta_{W}-\frac{1}{2}\right)\right]^{2}\left(\lambda_{ee}\lambda_{e\mu}m_{e}+\lambda_{e\mu}\lambda_{\mu\mu}m_{\mu}\right)^{2},$ (2)
where $\alpha$ is the fine structure constant, $G_{F}$ is the fermi constant, $\theta_{W}$ is the weak mixing angle, $M_{Z}$, $M_{Z^{\prime}}$, $m_{\mu}$ and $m_{e}$ is the mass of $Z^{0}$, $Z^{\prime}$, $\mu$ and $e$, respectively.
Since the contribution of $\lambda_{e\mu}\times\lambda_{ee}$ is suppressed by the ratio of lepton masses, we can assume that this process is entirely composed of $\lambda_{e\mu}\times\lambda_{\mu\mu}$ and is comparable with our result as shown in Fig. 7. Meanwhile, if considering the rigorous contributions, it set the joint limitations on these two coupling strength, as shown in Fig. 8.
@@
Figure 7: Sensitivity of the coupling strength $\lambda_{\mu\mu}\times\lambda_{e\mu}$ from $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ channel. Current and prospective results of $\mu\to e\gamma$ are include for comparison, assuming that it is entirely composed of $\lambda_{e\mu}\times\lambda_{\mu\mu}$.
Figure 8: Sensitivity of the combined coupling strength at the resonance point of the 1.55 TeV $e\mu$ collider. The results of $\mu\to eee$ and $\mu\to e\gamma$ are calculated by the current limits. Each shaded area serves as the exclusion zone for the corresponding process.